# SPIQA: A Dataset for Multimodal Question Answering on Scientific Papers

### v1

@@ C B-F1 L3S C B-F1 L3S C B-F1 L3S C B-F1 L3S C B-F1 L3S C B-F1 L3S
5.2 Main Results
@@
5.3 Ablation Study
Figure and Table Captions. In scientific papers, figures and tables are often accompanied by detailed and informative captions. Figure 3 shows that all Gemini and GPT variants experience significant performance drops when captions are excluded. For instance, the best-performing GPT-4o model suffers a 2-point L3Score decrease on test-A when captions are omitted, which underscores the importance of captions in providing context and enhancing comprehension in QA tasks.
Performance on Tables. Table 5 shows the performance of baseline models on tables and figures in test-A. Open-source models struggle with tables, scoring lower than their overall performance. Among closed-source models, Gemini 1.5 Pro, GPT-4 Vision, and GPT-4o perform well. Notably, GPT-4o achieves a 65.18 L3Score on tables, outperforming GPT-4 Vision, the second best by 8.50 points.
Performance on Figure Subcategories. Baseline models perform well on schematic diagrams but struggle with plots, charts, and visualizations. For example, GPT-4o scores 71.38 on schematics but only 56.73 and 56.07 on plots & charts, and visualizations, respectively. Such results indicate the need for improved systems to comprehend scientific diagrams requiring mathematical reasoning.
5.4 Qualitative Results and Error Analysis
Figure 4 illustrates two example questions, ground-truth answers, and outputs from various baseline models. For the plot-based question, both InstructBLIP and Gemini 1.5 Flash produce correct responses. However, only our proposed L3Score evaluates them correctly in both cases. Notably, InstructBLIP’s response (‘Ours’) correctly answers the question despite not matching the ground-truth. Traditional metrics like ROUGE-L and BERTScore fail to evaluate such responses, scoring them as 0. For the schematic-based question, the ground-truth is long and descriptive. GPT-4 Vision generates a significantly correct response, and L3Score accurately evaluates it with a score of 96.9. We observe that the best-performing GPT-4o struggles with complex plots, charts, and tables. Such error cases are detailed in the supplementary materials.
@@ 6 Conclusion
We introduce SPIQA (Scientific Paper Image Question Answering), the first large-scale QA dataset specifically designed to interpret complex figures and tables within the context of scientific papers. Additionally, we propose LLMLogScore, an improved metric for free-form QA that accurately analyzes the semantic context of candidate answers relative to the ground truth. Through extensive experiments with 12 prominent foundational models, we evaluate their ability to comprehend the nuanced aspects of research articles. Furthermore, fine-tuning two open-source systems, LLaVA and InstructBLIP, on the SPIQA training set results in significant improvements over zero-shot evaluations, indicating promising avenues for designing specialized systems for scientific QA in the future. Our work lays the foundation for developing advanced QA systems that effectively understand and analyze scientific documents, driving further research in this critical area.
@@ Appendix A Dataset and Code Release
Following NeurIPS Dataset and Benchmark Track guidelines, we publicly release the SPIQA dataset on Hugging Face: https://huggingface.co/datasets/google/spiqa. Our evaluation and metric computation scripts, along with responses from all closed- and open-source models, are accessible in the GitHub repository: https://github.com/google/spiqa. SPIQA will be properly maintained and will publicly remain available under the Creative Commons Attribution License (CC BY 4.0). The evaluation code and library for L3Score are available in the Github repository under Apache 2.0 license. We include the dataset card and README for the resources.
Appendix B Additional Ablation Study
@@
Full Text Figs & Tabs
BioAsq [70, 29] Human experts 3.2K -- PubMed Biomedical
BioRead [56] Cloze-style 16.4M 3.4M papers PubMed Biomedical
BioMRC [57] Cloze-style 812K 25M abstracts Pubtator Biomedical
emrQA [54] Cloze-style 455K 2.4K clinical notes i2b2 datasets EMRs
PubMedQA [20] Human experts 1K 120K abstracts PubMed Biomedical
BioASQ-QA [30] Human experts 4.7K -- PubMed Biomedical
ArgSciChat [63] Human Experts 41 dialogues 20 papers arXiv NLP
QASPER [14] Human experts 5K 1.5K papers S2ORC NLP
QASA [34] Human experts 1.8K 112 papers S2ORC AI/ML
Table C.1: Comparison of SPIQA with existing scientific question answering datasets. SPIQA is the first large-scale free-form QA corpus that focuses on the holistic aspects of scientific papers, including figures, tables, and full paper text. A dash (--) indicates that the number of papers or abstracts used in the final dataset is not reported.
@@
Closest to our work, QASA [34] contains 1,798 open-ended and detailed questions on 112 AI/ML papers, where annotators read the full paper text when writing the questions. The answers to these questions are multi-faceted and long-form, written by AI/ML experts or the actual authors of the corresponding papers. Due to the expensive annotation scheme, despite maintaining high quality, the QASA benchmark is small-scale and the questions do not require understanding complex figures and diagrams. Our proposed SPIQA dataset differs from existing scientific QA benchmarks in three key aspects: $(i)$ We introduce the first large-scale, free-form, open-ended scientific QA dataset covering all domains of computer science. $(ii)$ The questions and answers in SPIQA require a holistic understanding of complex figures and tables, along with the full textual content of the papers. $(iii)$ In addition to the direct QA setup, we propose a novel Chain-of-Thought (CoT) QA paradigm, where models first identify helpful figures and tables, and then generate the answer. This step-by-step QA pipeline helps evaluate the fine-grained reasoning capabilities of the baseline systems. Table C.1 provides an extensive comparison of SPIQA with all available scientific QA benchmarks.
@@ Appendix G Detailed Annotation Guidelines and User Interfaces
    Answering the question would require a complete understanding of the figure and its importance in the paper.

### v3

@@ C B-F1 L3S C B-F1 L3S C B-F1 L3S C B-F1 L3S C B-F1 L3S C B-F1 L3S
Method     CoT QA on SPIQA test-A
Ret. Acc. M R-L C B-F1 L3S
Human Eval 94.89 28.2 44.4 130.9 69.90 68.10
Gemini 1.5 Flash [63] 87.14 26.0 39.1 123.2 67.10 58.72
Gemini 1.5 Pro [63] 86.58 26.2 39.0 103.4 66.75 65.16
GPT-4 Vision [1] 84.30 26.0 38.6 110.1 66.94 64.55
GPT-4o [54] 86.58 27.4 43.7 130.2 68.93 66.36
Table 7: Human evaluation on a 20% subset of test-A on CoT QA. We also report the scores of baseline models on the same subset of test-A to contextualize their performance compared to human capability.
5.3 Ablation studies and performance on sub-categories
Captions are helpful. In scientific papers, figures and tables are often accompanied by detailed and informative captions. Figure 3 shows that all Gemini and GPT variants experience significant performance drops when captions are excluded. For instance, the best-performing GPT-4o model suffers a 2-point L3Score decrease on test-A when captions are omitted, which underscores the importance of captions in providing context and enhancing comprehension in QA tasks.
Figure Subcategories. Table 6 shows the performance of baseline models on tables and figures in test-A. Baseline models perform well on schematic diagrams but struggle with plots, charts, and visualizations. For example, GPT-4o scores 71.38 on schematics but only 56.73 and 56.07 on plots & charts, and visualizations, respectively. Such results indicate the need for improved systems to comprehend scientific diagrams requiring mathematical reasoning.
Tables harder for open models. Table 6 indicates open-source models struggle with tables, scoring lower than their overall performance. All closed-source models perform well. Notably, GPT-4o achieves a 65.18 L3Score on tables, outperforming GPT-4 Vision, the second best by 8.50 points.
Image resolution matters. Appendix Figure B.1 includes performance of the closed models on images of different resolutions. Full resolution images result in consistently better performance.
@@ 5.3 Ablation studies and performance on sub-categories
5.4 Human Evaluation
We conduct a human evaluation on a 20% subset of test-A to establish a human performance baseline for the CoT QA task in Table 7. The diversity of research domains in the SPIQA dataset poses a challenge in obtaining a reliable human upper bound, as it requires broad expertise to answer questions accurately. Additionally, the free-form response format yields varied answers depending on the evaluator’s knowledge and understanding of the figures and tables. To study these, we employ an experienced AI/ML researcher to evaluate a fraction of test-A, using the same setup as the large multimodal models. The human evaluator identifies relevant figures and tables and provides a free-form answer to the question, achieving a top-1 accuracy of 94.89% in retrieving the correct figure and outperforming the best model by approximately 8%. Unsurprisingly, on other automated metrics, the performance of human and LLMs are similar due to a single ground truth answer. To establish a more robust human baseline, future work should involve multiple human evaluators with diverse expertise.
5.5 Qualitative Results and Additional Analysis
Figure 4 illustrates two example questions, ground-truth answers, and outputs from various baseline models. For the plot-based question, both InstructBLIP and Gemini 1.5 Flash produce correct responses. However, only our proposed L3Score evaluates them correctly in both cases. Notably, InstructBLIP’s response (‘Ours’) correctly answers the question despite not matching the ground-truth. Traditional metrics like ROUGE-L and BERTScore fail to evaluate such responses, scoring them as 0. For the schematic-based question, the ground-truth is long and descriptive. GPT-4 Vision generates a significantly correct response, and L3Score accurately evaluates it with a score of 96.9. We observe that the best-performing GPT-4o struggles with complex plots, charts, and tables. More such error cases are detailed in Appendix H.
Spearman’s $\rho$ M R-L C B-F1
w. L3Score 0.71 0.72 0.69 0.78
Table 8: Correlation of L3Score with existing metrics for free-form QA evaluation.
Analysis of L3Score. Table 8 shows that the proposed L3Score metric correlates well with existing automated metrics. Appendix B presents additional analysis of the proposed metric. E.g., we compute L3Score using different LLMs (GPT 3.5, 4o, Gemini) and find it to be consistent when comparing the outputs of different baselines although the absolute score itself may be different depending on the model used for computing the L3Score.
@@ 6 Conclusion
We introduce SPIQA (Scientific Paper Image Question Answering), the first large-scale QA dataset specifically designed to interpret complex figures and tables within the context of scientific papers. Additionally, we propose LLMLogScore(L3Score), an improved metric for free-form QA that accurately analyzes the semantic context of candidate answers relative to the ground truth. Through extensive experiments with 12 prominent foundational models, we evaluate their ability to comprehend the nuanced aspects of research articles. Furthermore, fine-tuning two open-source systems, LLaVA and InstructBLIP, on the SPIQA training set results in significant improvements over zero-shot evaluations, indicating promising avenues for designing specialized systems for scientific QA in the future. Our work lays the foundation for developing advanced QA systems that effectively understand and analyze scientific documents, driving further research in this critical area.
@@ Appendix A Dataset and Code Release
Following NeurIPS Dataset and Benchmark Track guidelines, we publicly release the SPIQA dataset on Hugging Face: https://huggingface.co/datasets/google/spiqa. Our evaluation and metric computation scripts, along with responses from all closed- and open-source models, are accessible in the GitHub repository: https://github.com/google/spiqa. SPIQA is available under the Creative Commons Attribution License (CC BY 4.0). The evaluation code and library for L3Score are available in the Github repository under Apache 2.0 license. We include the dataset card and README for the resources.
Appendix B Additional studies on L3Score and ablations.
@@
Method Mode of Evaluation  Direct QA on SPIQA test-A
M R-L C B-F1 L3S
LLaVA 1.5 7B [40] Zero-shot 22.6 34.7 117.8 61.61 13.86
Unsupervised 22.6 35.1 118.2 61.53 14.43
Supervised 23.8 36.0 121.2 63.74 45.45
Unsupervised + Supervised 24.0 36.4 121.7 63.92 46.11
Table B.2: Ablation on fine-tuning strategy. Supervised fine-tuning yields significantly more improvements, showing the usefulness of the model-generated QAs on SPIQA training set.
Supervised vs. Unsupervised Tuning on SPIQA Training Set. To further understand the quality of model-generated QAs, we conduct an ablation study to investigate the impact of unsupervised training using figure-caption pairs from the SPIQA training set and report the results in Table B.2. During unsupervised training, we ask the model to describe a given figure, using the caption as ground truth. This approach does not require the generated QAs. As shown in the first two rows of Table B.2, such unsupervised training yields only marginal improvement over zero-shot evaluation on test-A. In contrast, supervised fine-tuning using QAs significantly enhances performance, demonstrating the value of generated annotations for advanced question-answering. Furthermore, we found that combining unsupervised figure-caption pre-training followed by QA-based fine-tuning leads to additional improvements, highlighting the potential benefits of large-scale QA data for developing specialized QA systems for scientific papers in the future.
@@
Closest to our work, QASA [35] contains 1,798 open-ended and detailed questions on 112 AI/ML papers, where annotators read the full paper text when writing the questions. The answers to these questions are multi-faceted and long-form, written by AI/ML experts or the actual authors of the corresponding papers. Due to the expensive annotation scheme, despite maintaining high quality, the QASA benchmark is small-scale and the questions do not require understanding complex figures and diagrams. Our proposed SPIQA dataset differs from existing scientific QA benchmarks in three key aspects: $(i)$ We introduce the first large-scale, free-form, open-ended scientific QA dataset covering all domains of computer science. $(ii)$ The questions and answers in SPIQA require a holistic understanding of complex figures and tables, along with the full textual content of the papers. $(iii)$ In addition to the direct QA setup, we propose a novel Chain-of-Thought (CoT) QA paradigm, where models first identify helpful figures and tables, and then generate the answer. This step-by-step QA pipeline helps evaluate the fine-grained reasoning capabilities of the baseline systems. Table 1 provides an extensive comparison of SPIQA with all available scientific QA benchmarks.
@@ Appendix G Detailed Annotation Guidelines and User Interfaces
    Answering the question would require understanding of the figure or table and may require domain knowledge.